# Statistical Guarantees of Correctness Coverage for Medical Multiple-Choice Question Answering

### v1

@@ Abstract
Large language models (LLMs) are increasingly deployed in real-world question-answering (QA) applications. However, LLMs have been proven to generate hallucinations and nonfactual information, undermining their trustworthiness in high-stakes medical tasks. Conformal prediction (CP) is well-known to be model-agnostic and distribution-free, which creates statistically rigorous prediction sets in classification tasks. In this work, we for the first time adapt the CP framework to medical multiple-choice question-answering (MCQA) tasks, by correlating the nonconformity score with the frequency score of correct options grounded in self-consistency theory, assuming no access to internal model information. Considering that the adapted CP framework can only control the (mis)coverage rate, we employ a risk control framework, which can manage task-specific metrics by devising a monotonically decreasing loss function. We evaluate our framework on 3 popular medical MCQA datasets utilizing 4 “off-the-shelf” LLMs. Empirical results demonstrate that we achieve user-specified average (or marginal) error rates on the test set. Furthermore, we observe that the average prediction set size (APSS) on the test set decreases as the risk level increases, which concludes a promising evaluation metric for the uncertainty of LLMs.
keywords:
Large Language Models, Conformal Prediction, Medical Multiple-Choice Question-Answering, Average Prediction Set Size
[…]
[inst1]organization=College of Electronic and Information Engineering, Tongji University, city=Shanghai, postcode=201804, country=China
@@ 1 Introduction
Recently, large language models (LLMs)[36] have gained prominence as transformative tools in real-world medical question-answering (QA) applications[27], offering significant potential to support both patients and healthcare professionals in effectively addressing medical inquiries [28, 4, 12]. In contemporary healthcare practices, a growing number of individuals rely on online search engines to obtain health-related information [13]. Despite the promising capabilities of LLMs, concerns regarding the reliability of their generated responses persist [7, 16], particularly in light of the potential hazards associated with inaccurate or contradictory information. Such risks are especially critical in clinical environments, where misinformation can lead to adverse outcomes [24].
@@
Uncertainty quantification (UQ) has emerged as a critical methodology to assess the reliability of LLM outputs [20, 19, 33, 6, 26]. Existing UQ techniques can be broadly categorized into white-box and black-box approaches. White-box methods typically utilize internal model information, such as logits, to quantify semantic uncertainty [21, 9, 32], including metrics like semantic entropy [10]. In contrast, black-box uncertainty measures, rooted in self-consistency theory, evaluate uncertainty based on the semantic diversity of multiple outputs generated for the same input, without requiring access to the model’s internal states [22, 31, 25]. Despite their utility, existing methods rely on heuristic definitions of uncertainty and lack rigorous statistical guarantees for task-specific performance metrics [18, 34].
Figure 1: Overview of our framework.
@@
Conformal prediction (CP) has emerged as a robust framework within the broader domain of machine learning, offering statistically rigorous guarantees of marginal (average) coverage for prediction sets [2, 5]. Previous research has demonstrated the successful adaptation of CP to classification tasks [3, 1, 15], achieving user-defined correctness miscoverage rates. However, extending CP to medical question-answering (QA) tasks presents substantial challenges. In this work, we introduce a novel adaptation of the CP framework, which develops the nonconformity score (NS) to be closely aligned with the correct option in medical multiple-choice question-answering (MCQA) tasks. Specifically, we define the NS as one minus the probability of the correct option for each calibration data point. Given that access to internal model information may not be feasible in practical medical MCQA scenarios, we generate multiple outputs for the same medical query and compute the frequency score of each option. Leveraging self-consistency theory [30], we utilize the frequency score within the sampling set as a proxy for the option’s probability. Furthermore, Given that our adapted framework only provides rigorous control over the coverage of correct options, we further present a risk control framework to manage task-specific metrics in Section 3. When we devise the loss function to reflect the correctness miscoverage, we can achieve equivalent results.
@@
We evaluate our approach on 3 multiple medical MCQA datasets, including the MedMCQA [17], MedQA [17], and MMLU [14] datasets, utilizing 4 popular pre-trained LLMs, including Llama-3.2-1B-Instruct, Llama-3.2-3B-Instruct, Qwen2.5-1.5B-Instruct, and Qwen2.5-3B-Instruct. Empirical results demonstrate that we achieve rigorous management of the marginal error rate (average correctness miscoverage rate) on the fresh test MCQA samples at various risk levels. Moreover, we also estimate the average prediction set size (APSS) within the test set. Experimental findings show that the APSS metric decreases as the risk level increases, which can assist in using indicators such as accuracy to comprehensively evaluate the performance of LLMs. Furthermore, We conduct comprehensive ablation studies, which illustrate the robustness and effectiveness of our framework. Our contributions can be summarized as follows:
  • 1.
    We, for the first time, adapt conformal prediction to medical multiple-choice question-answering tasks, which rigorously guarantees the coverage of correct options.
  • 2.
    We disclose a promising metric to evaluate the uncertainty of LLMs by estimating the average prediction set size on test MCQA samples.
  • 3.
    We conduct extensive evaluations and ablation studies on 3 popular medical benchmarks utilizing 4 pretrained LLMs, demonstrating the effectiveness and robustness of our framework.
@@ 2 Related Work
2.1 NLG tasks in the medical domain
The application of Natural Language Generation (NLG) in the medical domain has become a critical area of research, driven by the growing reliance on advanced technologies to optimize patient care and clinical communication. NLG systems are increasingly employed to automate clinical documentation, address medical inquiries, and provide diagnostic suggestions. These systems utilize state-of-the-art large language models (LLMs) such as GPT-3 [4] and BERT [8] to produce responses that are both contextually relevant and medically accurate. By delivering timely and precise information, these technologies aim to alleviate the workload of healthcare professionals and contribute to enhanced patient outcomes [24].
While NLG models present considerable potential within medical applications, they also encounter substantial challenges. A primary concern is the reliability of the generated content, especially in high-stakes scenarios such as medical diagnosis and patient care. Despite their ability to generate contextually appropriate and persuasive responses, these models may also produce factually incorrect or misleading information, posing significant risks in clinical environments. Recent research efforts have focused on implementing domain-specific fine-tuning strategies and integrating uncertainty quantification (UQ) techniques to assess and enhance the reliability of NLG systems in medical contexts [22]. Nonetheless, further advancements are needed to ensure that NLG systems consistently meet the rigorous safety and accuracy standards required in healthcare settings.
2.2 UQ in medical QA tasks
@@
Recent advancements in the field have introduced specialized UQ techniques tailored to the unique demands of medical QA tasks. Confidence-based metrics [20] and entropy-based approaches [23] have been employed to evaluate the uncertainty associated with medical responses. These methods typically involve analyzing the distribution of model outputs or assessing the consistency of predictions across multiple generated responses. Additionally, self-consistency frameworks and ensemble methods have been proposed to compute confidence scores by aggregating the outputs from diverse model iterations [22]. Despite their potential, many existing approaches lack formal statistical guarantees, representing a significant limitation in clinical settings where precision and reliability are critical. Consequently, there remains a pressing need for the development of robust UQ methodologies capable of effectively managing the complexities of medical data while delivering reliable and interpretable uncertainty estimates.
2.3 CP
Conformal Prediction (CP) has emerged as a robust methodology for uncertainty quantification (UQ) in medical question-answering (QA) tasks, providing a rigorous framework for constructing prediction sets with statistically guaranteed coverage [5]. Unlike traditional UQ techniques that yield single-point predictions, CP generates a set of potential outcomes alongside a confidence measure, offering a quantifiable assessment of the reliability of a generated response. This capability is particularly critical in medical QA applications, where the implications of erroneous predictions can be severe, and where clinicians must often rely on automated responses to inform clinical decision-making.
The utility of CP is further enhanced in the context of large language models (LLMs)[35, 29, 31, 11, 1], which are frequently characterized as ”black-box” models due to their inherent opacity [22]. A key advantage of CP is its model-agnostic nature, as it does not require assumptions about the underlying data distribution, allowing its application across a broad spectrum of machine learning models, including LLMs. Recent studies have explored the adaptation of CP methodologies to open-ended natural language generation (NLG) tasks, such as medical QA, demonstrating its potential to balance prediction accuracy with coverage. This controlled trade-off is particularly valuable in high-stakes domains like healthcare, where maintaining a balance between model performance and reliability is essential. Nevertheless, significant challenges remain in scaling CP approaches to accommodate large models and managing the complexities associated with diverse and heterogeneous medical data.
@@ 3.1 Adaptation of Conformal Prediction to Medical MCQA Tasks
Conformal Prediction (CP) is a statistical framework that generates prediction sets with robust uncertainty quantification for machine learning models. Unlike traditional methods that provide single-point predictions, CP constructs a set of potential outcomes for each test instance, ensuring with a confidence level of at least $1-\alpha$ that the true label is included. This characteristic is particularly valuable for high-stakes applications such as medical MCQA tasks.
In the context of medical MCQA, CP calculates a conformity score $s(X_{\text{test}},Y_{\text{test}})$ for each test instance, comparing it with the empirical distribution of conformity scores from the training data. A threshold $\hat{q}$ is then derived to define the prediction set $C(X_{\text{test}})$ that meets the desired coverage probability:
@@
The calibration process further enhances this reliability by fine-tuning the threshold using a dedicated calibration set. However, while CP provides a robust approach to managing uncertainty in predictive models and offers well-calibrated prediction sets suitable for the stringent requirements of medical MCQA tasks, its limitation lies in the inability to manage task-specific metrics beyond mere coverage control. This shortcoming underscores the necessity of our proposed risk control framework, which extends the capabilities of CP by introducing a monotonically decreasing loss function to effectively manage task-specific risks and enhance the reliability of large language models in high-stakes medical applications.
@@ 3.2 Guarantee of Task-specific Metrics
In this section, we present a method for constructing prediction sets with statistically rigorous miscoverage rate guarantees through conformal calibration. The proposed approach leverages a calibration set $\{(x_{i},y_{i}^{*})\}_{i=1}^{n}$ and a test data point $(x_{n+1})$ to achieve precise task-specific risk management through the design of tailored loss functions.
The proposed method extends traditional classification tasks into the Question Answering (QA) domain by leveraging the flexibility of the Non-Conformity Score. In a typical classification task, a model predicts a discrete label from a predefined set of classes. By mapping each class label to a potential answer in a natural language format, the classification task is transformed into a QA task. This mapping is particularly effective in scenarios where the model generates candidate answers that require more contextual understanding and natural language generation.
For example, a classification task that categorizes images into ”cat,” ”dog,” or ”bird” can be rephrased as a QA task by asking, ”What animal is in the image?” The candidate answers become the possible labels, and the model evaluates each candidate’s reliability through the Non-Conformity Score, ensuring that the prediction set adheres to the desired coverage guarantees.
@@
where $R(y)$ is a quantitative measure of the reliability of a candidate answer $y$ within the set. The parameter $t$ functions as a threshold, systematically regulating the inclusion of predictions based on their reliability scores.
@@
where $R(y)$ represents the reliability score of the candidate answer. The Non-Conformity Score serves as a critical component in bridging classification and QA tasks by providing a unified metric for evaluating prediction quality across different task types.
@@
where $y_{i}^{*}$ is the ground-truth answer. The loss function is constructed to be non-increasing with respect to the tuning parameter $t$, ensuring that a higher parameter value does not lead to an increased loss. This property is critical for maintaining the theoretical monotonicity required for robust empirical risk management.
@@
The primary objective of the proposed method is to control the expected loss for new test data by ensuring:
@@
This formulation guarantees a statistically rigorous control of the miscoverage rate, thereby offering robust theoretical assurances for predictive accuracy and reliability. The method not only ensures a statistically rigorous miscoverage rate but also provides task-specific metric management through customized loss function design. By tailoring the loss function to meet the specific requirements of both classification and QA tasks, the proposed approach offers a highly adaptable risk management strategy. The integration of the Non-Conformity Score further enhances the method’s ability to maintain consistent risk control across varying prediction scenarios, highlighting its potential in high-stakes applications requiring stringent statistical guarantees.

### v2

# Correctness Coverage Evaluation for Medical Multiple-Choice Question Answering Based on the Enhanced Conformal Prediction Framework

@@ Abstract
Large language models (LLMs) are increasingly adopted in medical question-answering (QA) scenarios. However, LLMs can generate hallucinations and nonfactual information, undermining their trustworthiness in high-stakes medical tasks. Conformal Prediction (CP) provides a statistically rigorous framework for marginal (average) coverage guarantees but has limited exploration in medical QA. This paper proposes an enhanced CP framework for medical multiple-choice question-answering (MCQA) tasks. By associating the non-conformance score with the frequency score of correct options and leveraging self-consistency, the framework addresses internal model opacity and incorporates a risk control strategy with a monotonic loss function. Evaluated on MedMCQA, MedQA, and MMLU datasets using four off-the-shelf LLMs, the proposed method meets specified error rate guarantees while reducing average prediction set size with increased risk level, offering a promising uncertainty evaluation metric for LLMs.
Keywords Large Language Models  $\cdot$ Conformal Prediction  $\cdot$ Medical Multiple-Choice Question-Answering  $\cdot$ Average Prediction Set Size
@@ 1 Introduction
Recently, large language models (LLMs)[1, 2, 3, 4, 5] have gained prominence as revolutionary instruments in real-world medical question-answering (QA) applications[6, 7], offering significant potential to help both patients and healthcare professionals effectively address medical inquiries [8, 9, 10]. In contemporary healthcare practices, many people rely on online search engines to obtain health-related information [11]. Despite the promising capabilities of LLMs, concerns persist about the reliability of their generated responses [12, 13, 14], particularly in light of the potential hazards associated with inaccurate or contradictory information. Such risks are especially critical in clinical settings, where misinformation can lead to adverse outcomes and potentially endanger patient safety [15]. Unlike general-domain QA tasks, medical QA often involves domain-specific terminology, context-sensitive reasoning, and high-stakes decision-making, where even minor inaccuracies may have significant consequences. Therefore, ensuring the reliability and trustworthiness of model outputs is not merely a performance issue, but a fundamental requirement for real-world applicability. These challenges highlight the need for rigorous uncertainty quantification frameworks with formal guarantees, and strongly motivate the adaptation and extension of conformal prediction methods specifically for the medical QA domain.
@@
Uncertainty quantification (UQ) has emerged as a critical methodology for assessing the reliability of LLM output [16, 17, 18, 19, 20]. Existing UQ techniques can be broadly categorized into white-box and black-box approaches. White-box methods utilize internal model information, such as logarithmic summation, to quantify semantic uncertainty [21, 22, 23], including metrics such as semantic entropy [24]. In contrast, black-box uncertainty measures, rooted in self-consistency theory, evaluate uncertainty based on the semantic diversity of multiple outputs generated for the same input without requiring access to the model’s internal states [25, 26, 27]. Despite their utility, existing methods rely on heuristic definitions of uncertainty and lack of rigorous statistical guarantees for task-specific performance metrics [28].
Conformal prediction (CP) is now recognized as a robust framework within the broader domain of machine learning, offering statistically rigorous guarantees of marginal (average) coverage for prediction sets [29, 30]. Previous research has demonstrated the successful adaptation of CP to classification tasks [31, 32, 33], achieving error correction miscoverage rates defined by the user. However, extending CP to medical question-answering (QA) tasks presents substantial challenges.
@@
To address these limitations, this study proposed an enhanced CP framework, which develops the non-conformity score (NS) to be closely aligned with the correct option in multiple-choice medical question-answering (MCQA) tasks. Given that access to internal model information may not be feasible in practical medical MCQA scenarios, the enhanced CP framework defines the NS as one minus the probability of the correct option for each calibration data point, generates multiple outputs for the same medical query, and computes the frequency score for each option. Using self-consistency theory [34], the enhanced CP framework utilizes the frequency score within the sampling set as a proxy for the probability of the option. Furthermore, to address the limitation that the enhanced CP framework only provides rigorous control over the correct answer coverage, this study presents a risk control framework to manage task-specific metrics, ensuring that equivalent results can be obtained when the loss function is designed to reflect the correctness error miscoverage. So that equivalent results can be obtained when the loss function is designed to reflect the correctness error miscoverage
@@
The proposed enhanced CP framework was evaluated on three multiple medical MCQA datasets, including the MedMCQA [35], MedQA [35], and MMLU [36] datasets, utilizing four popular pre-trained LLMs, including Llama-3.2-1B-Instruct, Llama-3.2-3B-Instruct, Qwen2.5-1.5B-Instruct, and Qwen2.5-3B-Instruct. Empirical results demonstrate that the enhanced CP framework achieves strict control over the marginal error rate (average correctness miscoverage rate) on test MCQA samples at various risk levels. Moreover, the average prediction set size (APSS) within the test set is also estimated in this study. Experimental findings show that the APSS metric decreases as the risk level increases, which can assist in using indicators such as accuracy to comprehensively evaluate the performance of LLMs. The main contributions of this study are summarized as follows. (i) For the first time, conformal prediction is applied to the medical multiple-choice question-answering tasks, strictly guaranteeing the coverage of correct options. (ii) A promising metric is revealed to assess the uncertainty of LLM by estimating the average prediction set size on the test MCQA sample. (iii) Extensive evaluation and ablation studies on three popular medical benchmarks are conducted using four pre-trained LLMs, demonstrating the effectiveness and robustness of the enhanced CP framework.
The remainder of this study is organized as follows. Section 2 introduces related works. In section 3, a detailed enhanced CP framework has been illustrated and explained. An experiment is presented in section 4. Section 5 concludes this study and gives future works.
@@ 2 Related Work
Medical question-answering (QA) presents distinct challenges compared to general-domain QA tasks, owing to its reliance on domain-specific terminology, the necessity for precise multi-hop reasoning across heterogeneous medical knowledge sources, and the potentially severe consequences of erroneous outputs. Furthermore, many medical queries are inherently ambiguous or highly context-dependent, and the corresponding data are often incomplete or uncertain.
These complexities hinder the ability of conventional QA models to produce reliable responses, particularly in open-ended or generative settings. To mitigate these limitations, recent research has introduced self-consistency-based approaches, which enhance model robustness by measuring the agreement across multiple independently generated responses. Building upon this paradigm, the present study integrates self-consistency mechanisms within a conformal prediction framework, enabling the construction of uncertainty-aware prediction sets that offer formal statistical coverage guarantees, while maintaining task-specific reliability in high-stakes medical QA scenarios.
2.1 Natural Language Generation tasks in the medical domain
The application of Natural Language Generation (NLG) in the medical domain has become a critical area of research, driven by the growing reliance on advanced technologies to optimize patient care and clinical communication. NLG systems are increasingly employed to automate clinical documentation, address medical inquiries, and provide diagnostic suggestions. These systems utilize state-of-the-art large language models (LLMs) such as GPT-3 [9] and BERT [37] to produce responses that are both contextually relevant and medically accurate. By delivering timely and precise information, these technologies aim to alleviate healthcare professionals’ workload and enhance patient outcomes [15].
While NLG models present considerable potential within medical applications, they also encounter substantial challenges. A primary concern is the reliability of the generated content, especially in high-stakes scenarios such as medical diagnosis and patient care. Despite their ability to develop contextually appropriate and persuasive responses, these models may also produce factually incorrect or misleading information, posing significant risks in clinical environments. Recent research efforts have focused on implementing domain-specific fine-tuning strategies and integrating uncertainty quantification (UQ) techniques to assess and enhance the reliability of NLG systems in medical contexts [25]. Nonetheless, further advancements are needed to ensure that NLG systems consistently meet the rigorous safety and accuracy standards of healthcare settings.
2.2 Uncertainty Quantification in Medical Question-Answering Tasks
@@
Recent advancements in the field have introduced specialized UQ techniques tailored to the unique demands of medical QA tasks. Confidence-based metrics [16] and entropy-based approaches [38] have been employed to evaluate the uncertainty associated with medical responses. These methods typically involve analyzing the model’s output distribution or assessing the consistency of predictions across multiple generated responses. Additionally, self-consistency frameworks and ensemble methods have been proposed to compute confidence scores by aggregating the outputs from diverse model iterations [25]. Despite their potential, many existing approaches lack formal statistical guarantees, representing a significant limitation in clinical settings where precision and reliability are critical. Consequently, developing robust UQ methodologies is necessary to effectively manage the complexities of medical data while delivering reliable and interpretable uncertainty estimates.
2.3 Conformal Prediction for Uncertainty Quantification in Medical Question-Answering Tasks
Conformal Prediction (CP) has emerged as a robust methodology for uncertainty quantification (UQ) in medical question-answering (QA) tasks, providing a rigorous framework for constructing prediction sets with statistically guaranteed coverage [30, 39]. Unlike traditional UQ techniques that produce single-point predictions, CP generates a set of potential outcomes alongside confidence measures, offering a quantifiable assessment of the reliability of a generated response. The capability is particularly critical in medical QA applications, where the implications of erroneous predictions can be severe, and clinicians must often rely on automated responses to inform clinical decision-making.
The utility of CP is further enhanced in the context of large language models (LLMs)[40, 41, 26, 42, 32], which are frequently characterized as “black-box” models due to their inherent opacity [25]. A key advantage of CP is its model-agnostic nature, as it does not require assumptions about the underlying data distribution, allowing its application across a broad spectrum of machine learning models, including LLMs. Recent studies have explored the adaptation of CP methodologies to open-ended natural language generation (NLG) tasks, such as medical QA, demonstrating the potential of CP to balance prediction accuracy with coverage. This controlled trade-off is particularly valuable in high-stakes domains like healthcare, where maintaining a balance between model performance and reliability is essential. Nevertheless, significant challenges remain in scaling CP approaches and managing the complexities of diverse and heterogeneous medical data.
@@ 3 Method
The methods presented in this section introduce a novel framework based on Conformal Prediction (CP) for medical multiple-choice question-answering tasks. This framework provides an advanced approach for uncertainty quantification (UQ) and task-specific risk management. By leveraging the framework, prediction sets are generated that are statistically guaranteed to contain the true label with a predefined confidence level, which is essential for applications in high-stakes fields such as healthcare.
@@ 3.1 Adaptation of Conformal Prediction to Medical MCQA Tasks
Conformal Prediction (CP) is a statistical framework that generates prediction sets with robust uncertainty quantification for machine learning models. Unlike traditional methods that provide single-point predictions, CP constructs a set of potential outcomes for each test instance, ensuring that the true label is included with at least $1-\alpha$ confidence. This characteristic is particularly valuable for high-stakes applications such as medical MCQA tasks.
As illustrated in Figure 1, the framework first measures the disagreement between the input $x_{i}$ and its ground-truth label $y_{i}^{*}$ by computing a non-conformity score for each calibration sample. A quantile threshold $\hat{q}$ is then determined based on these scores to establish the statistical risk level. Given a test MCQA sample $x_{\text{test}}$, the model evaluates multiple candidate answers, computing their respective non-conformity scores. The prediction set is subsequently constructed by retaining only those answer choices whose scores fall below the calibrated threshold, ensuring that the final prediction set adheres to the desired miscoverage rate guarantees.
Figure 1: Adaptation of Conformal Prediction to Medical MCQA Tasks.
In medical MCQA, CP calculates a conformity score $s(X_{\text{test}},Y_{\text{test}})$ for each test instance, comparing it with the empirical distribution of conformity scores from the training data. A threshold $\hat{q}$ is then derived to define the prediction set $C(X_{\text{test}})$ that meets the desired coverage probability:
@@
The calibration process further enhances this reliability by fine-tuning the threshold using a dedicated calibration set. However, while CP provides a robust approach to managing uncertainty in predictive models and offers well-calibrated prediction sets suitable for the stringent requirements of medical MCQA tasks, its limitation lies in the inability to manage task-specific metrics beyond mere coverage control. This shortcoming underscores the necessity of the proposed risk control framework, which extends the capabilities of CP by introducing a monotonic decreasing loss function to manage task-specific risks effectively and enhance the reliability of large language models in high-stakes medical applications.
@@ 3.2 Guarantee of Task-specific Metrics
This section introduces a method for constructing prediction sets with statistically rigorous guarantees on miscoverage rates via conformal calibration, specifically tailored for the medical multiple-choice question answering (MCQA) context. Unlike general-purpose QA tasks, medical MCQA involves domain-specific, often rare and compound terminology, which is typically segmented into multiple sub-tokens during language model processing. This characteristic introduces inconsistency in token-level semantic evaluations, leading to unstable and biased uncertainty estimates. To address this challenge, the proposed framework operates at the answer option level, treating each candidate response as a discrete class and computing non-conformity scores to quantify its deviation from correctness. By leveraging a calibration set $\{(x_{i},y_{i}^{*})\}_{i=1}^{n}$ and a test data point $(x_{n+1})$ , the method enables precise task-specific risk control through the design of customized loss functions, facilitating robust uncertainty quantification in high-stakes medical scenarios.
The proposed method extends traditional classification tasks into the Question Answering (QA) domain by leveraging the flexibility of the Non-Conformity Score. In a typical classification task, a model predicts a discrete label from a predefined set of classes. The classification task is transformed into a QA task by mapping each class label to a potential answer in a natural language format. This mapping is particularly effective when the model generates candidate answers requiring more contextual understanding and natural language generation.
For example, a classification task that categorizes images into “cat,” “dog,” or “bird” can be rephrased as a QA task by asking, “What animal is in the image?” The candidate answers become the possible labels, and the model evaluates each candidate’s reliability through the Non-Conformity Score, ensuring that the prediction set adheres to the desired coverage guarantees.
@@
$R(y)$ is a quantitative measure of the reliability of a candidate’s answer $y$ within the set. The parameter $t$ functions as a threshold, systematically regulating the inclusion of predictions based on their reliability scores.
@@
$R(y)$ represents the reliability score of the candidate answer. The Non-Conformity Score is critical in bridging classification and QA tasks by providing a unified metric for evaluating prediction quality across different task types.
@@
$y_{i}^{*}$ is the ground-truth answer, the loss function is designed as a non-increasing function with respect to the tuning parameter $t$, ensuring that a higher parameter value does not increase the loss. This property is critical for maintaining the theoretical monotonicity required for robust empirical risk management.
@@
The primary objective of the proposed method is to control the expected loss of new test data by ensuring:
@@
This formulation guarantees a statistically rigorous control of the miscoverage rate, thereby offering robust theoretical assurances for predictive accuracy and reliability. Beyond ensuring marginal coverage, the proposed method provides flexible support for task-specific metric control through a customized loss function design. By tailoring the loss function to the characteristics of classification and question-answering (QA) tasks, this framework presents a highly adaptable risk management strategy. Integrating the non-conformity score further enhances the model’s ability to maintain stable uncertainty control across various prediction scenarios, which is particularly beneficial in high-stakes applications requiring strong statistical guarantees.
The complete procedure is summarized in Algorithm 1, which details each step in constructing prediction sets with conformal guarantees for medical MCQA tasks, and demonstrates how statistical rigor is operationalized into a practical inference framework.
Input: Calibrated sampled generations $\mathcal{D}^{c}=\left\{(x_{i},\{y_{j}^{i}\},y_{i}^{*})\right\}_{i=1}^{m}$;
Test sample $x_{m+1}$ with sampled responses $\{y_{j}^{(m+1)}\}$;
User-specified significance level $\varepsilon$; Class number $C$.
1
2foreach $i\in\{1,\dots,m\}$ do
3       Count the occurrences of each option $A$ to $D$ in $\{y_{j}^{i}\}$ to estimate empirical probabilities $\hat{p}_{i,1},\dots,\hat{p}_{i,C}$
4
5foreach $i\in\{1,\dots,m\}$ do
6       Let $y_{i}^{*}$ be the correct label for $x_{i}$
7       Compute non-conformity score: $\alpha_{i}=1-\hat{p}_{i,y_{i}^{*}}$
8
9Count class frequencies in $\{y_{j}^{(m+1)}\}$ to obtain $\hat{p}_{m+1,1},\dots,\hat{p}_{m+1,C}$
10for $j\leftarrow 1$ to $C$ do
11       $\alpha_{j}^{(m+1)}=1-\hat{p}_{m+1,j}$
12       $p\left(Y_{j}^{(m+1)}\right)=\frac{1}{m+1}\cdot\#\left\{i:\alpha_{i}\geq\alpha_{j}^{(m+1)}\right\}$
13       if $p\left(Y_{j}^{(m+1)}\right)>\varepsilon$ then
14             Add $Y_{j}^{(m+1)}$ to prediction set $\Gamma^{\varepsilon}$
15
16
17Prediction $\leftarrow\arg\max_{j}\hat{p}_{m+1,j}$
Output: Prediction, $\Gamma^{\varepsilon}$
Algorithm 1 Frequency-Based Implementation of Conformal Prediction for Medical MCQA